The floor of a one story building is 14 feet longer than is wide. The building has 1472 square feet of floor space:
a) write quadratic equation for the area of the floor in terms of width.
b) find the length and width of the floor:
w=
l=

Answer by Boreal(15235)

width is x
length is x+14
x*(x+14)=1472
x^2+14x-1472=0
That is the quadratic equation.
(x+46)(x-32)=0
x=-46 or 32; the first can't be possible
width is 32 feet
length is 46 feet
area is 32*46=1472 sq ft